### v1

# Nonmonotonic diffusion in sheared active suspensions of squirmers

Abstract
We investigate how shear influences the dynamics of active particles in dilute to concentrated suspensions. Specifically, we focus on apolar active suspensions of squirmers, which are individually immotile but display activity-induced hydrodynamic diffusion. Under shear, the instantaneous particle velocities fluctuate more rapidly, similar to passive suspensions; however, surprisingly, the long-time diffusive dynamics can slow down and exhibit a nonmonotonic dependence on the shear rate. We show that this behavior arises from an interplay between the activity-induced persistent motion and shear-induced negative velocity autocorrelation, both of which are more pronounced at lower volume fractions. Simulations of self-propelled particles with tunable persistence support this explanation and further offer a simple mechanism to understand the observed coupling. Our results reveal a generic effect of shear on diffusion in active suspensions, elucidate how internal and external forcing interact, and provide new possibilities to modulate transport in active fluids.
Suspending particles able to inject energy in a fluid provides a means to drive the fluid out of equilibrium internally. Such active suspensions, or active fluids, have been the subject of intense study in recent decades, showing rich phenomena unseen in their passive, equilibrium counterparts [1, 2]. As a typical example, bacterial suspensions have been found to display enhanced diffusion [3, 4, 5, 6, 7, 8] and turbulent-like chaotic motion [9, 10, 11, 12, 13] in the absence of any external flow, as well as upstream swimming [14, 15, 16] and superfluid-like rheology [17, 18, 19] under pressure or shear-driven flows. Understanding how various internal and external driving mechanisms interact across spatial and temporal scales is important for predicting and controlling the dynamics and mechanics of active fluids in complex conditions, yet a detailed knowledge of the interplay has not been established.
One of the simplest ways to drive a fluid externally is to shear it at a steady, homogeneous shear rate, $\dot{\gamma}$. When sheared, passive noncolloidal suspensions undergo shear-induced diffusion (SID) in the plane perpendicular to the mean flow [20, 21, 22, 23]. SID arises from fluid-mediated interparticle collisions, and dimensional analysis suggests that the resulting diffusion coefficient, $D_{s}$, scales as $\dot{\gamma}a^{2}f(\phi)$, where $a$ is the particle radius and $f(\phi)$ a function of the particle volume fraction, $\phi$ [24]. The precise form of $f(\phi)$ depends on the physical regime as well as the shape and surface properties of the particles [25, 26, 27, 28, 29]. However, regardless of the details, one expects $D_{s}$ to increase with $\dot{\gamma}$ because particles collide more frequently at higher $\dot{\gamma}$.
In this Letter, we investigate the interplay between shear and activity, and examine whether increasing shear rate enhances diffusion in active fluids. To this end, we use a recently modified Fast Stokesian Dynamics (FSD) method [30, 31] to perform large-scale hydrodynamic simulations of dilute to concentrated active suspensions ($\phi\in[10\%,40\%]$). Specifically, we simulate apolar active suspensions of “shakers,” particles that are active but not self-propelling, which collectively display activity-induced hydrodynamic diffusion [31]. The dynamics can be characterized by a diffusion coefficient, $D_{0}$, allowing for a Péclet number, Pe $\equiv\dot{\gamma}a^{2}/D_{0}$, to be defined to measure the relative effect of shear versus activity. Extensive simulations of thousands of shakers at different Pe and $\phi$ show that such suspensions always reach diffusive dynamics at long times; however, surprisingly, the effective diffusion coefficient, $D$, can drop below $D_{0}$ for a range of Pe when $\phi$ is small. For example, $D/D_{0}\approx 0.4$ when Pe $\approx 100$ and $\phi=10\%$. As Pe increases further, $D/D_{0}$ rises again, with shear eventually dominating, resulting in nonmonotonic diffusion. We analyze this behavior in detail and trace its physical origin to an activity-induced persistent motion and a shear-induced negative velocity autocorrelation, both of which are more pronounced at lower volume fractions. Additional simulations of self-propelled particles with tunable persistence support this interpretation and further provide a simple mechanism to understand the observed coupling. By contrast, the relative diffusivity of passive colloids with negligible persistence increases monotonically with Pe, as expected. Overall, our results demonstrate a general effect of shear on diffusion in active suspensions, elucidate how interactions between internal and external forcing influence these complex fluids, and suggest new ways to control mixing and mass transport in active fluids.
Figure 1: Nonmonotonic diffusion in sheared active suspensions. (a) Illustration of a typical simulation, showing 1024 shakers at 10% volume fraction under shear. (b) Mean square displacements in the $yz$-plane for pulling shakers or passive particles under different shear rates ($\phi=10\%$). The inset shows a few sample trajectories of shakers at $\dot{\gamma}=10^{-3}$ and 1. (c) Relative diffusion coefficient, $D/D_{0}$, against Pe for pulling or pushing shakers at different $\phi$.
Models and methods—We model active particles as “squirmers,” spheres with a prescribed axisymmetric slip velocity [32, 33, 34]. The surface slip at position $\bm{r}$ from the sphere’s center may be expressed as
$\bm{u}_{\text{s}}=-\bm{U}_{\text{s}}-\bm{\Omega}_{\text{s}}\times\bm{r}+\bm{r}\cdot\bm{E}_{\text{s}}+\dots,$ (1)
where $\bm{U}_{\text{s}}$ is the self-propulsion velocity, $\bm{\Omega}_{\text{s}}$ the self spin, $\bm{E}_{\text{s}}$ the self strain rate, and dots denote terms involving higher order irreducible tensors of $\bm{r}$ [35]. Specifically,
$\bm{U}_{\text{s}}=\frac{2}{3}B_{1}\bm{p},\ \bm{\Omega}_{\text{s}}=\frac{C_{1}}{a^{3}}\bm{p},\ \bm{E}_{\text{s}}=-\frac{3}{5}\frac{B_{2}}{a}\big{(}\bm{p}\bm{p}-\frac{1}{3}\bm{I}\big{)},$ (2)
where $B_{1}$ and $B_{2}$ are the leading-order squirming modes, $C_{1}$ a coefficient associated with azimuthal slip, $\bm{p}$ the particle director, and $\bm{I}$ an identity tensor [35]. Setting $B_{1}=0$ yields an immotile squirmer known as a shaker.
Although shakers are individually immotile, they swim collectively due to hydrodynamic interactions (HIs). To leading order, the HIs arise from $\bm{E}_{\text{s}}$ if $B_{2}\neq 0$: $B_{2}>0$ corresponds to pullers, while $B_{2}<0$ pushers. Therefore, the ratio $\beta\equiv B_{2}/B_{1}$ indicates the type of swimmer and the relative importance of HIs and self-propulsion. In this work, we mainly take $\beta\to\pm\infty$ to amplify the effect of HIs, but have also considered neutral squirmers with $\beta=0$ and additional rotational diffusion of $\bm{p}$ on each particle, effectively modeling active Brownian particles (ABPs, discussed later). Higher-order squirming modes and the azimuthal slip ($C_{1}$) are neglected, as is common in the literature [36, 37, 38, 31].
The dynamics of a suspension of squirmers is described by the active Stokesian Dynamics framework [35]. The governing equations can be written compactly as
$\mathbf{U}=\mathbf{U}_{\infty}+\mathbf{U}_{\text{s}}+\mathcal{R}_{\mathrm{FU}}^{-1}\bm{\cdot}\big{(}\mathcal{R}_{\text{FE}}\bm{:}(\mathbf{E}_{\infty}-\mathbf{E}_{\text{s}})+\mathbf{F}_{\text{ext}}\big{)},$ (3)
where the velocities ($\mathbf{U},\mathbf{U}_{\infty},\mathbf{U}_{\text{s}}$), strain rates ($\mathbf{E}_{\infty},\mathbf{E}_{\text{s}}$) and force ($\mathbf{F}_{\text{ext}}$) are all $N$-tuples, with elements defined at the particle centers, e.g., ${\mathbf{U}}\equiv(\bm{U}_{1}\ \bm{U}_{2}\ ...\ \bm{U}_{N})^{T}$, and ${\mathcal{R}}_{\text{FU}}$ and ${\mathcal{R}}_{\text{FE}}$ are the resistance tensors coupling the force and velocity moments of all $N$ particles. Note how individual particle activities couple to the large-scale flows through HIs. As illustrated in Fig. 1(a), we apply a shear in the $x$ direction, $\bm{U}_{\infty}(x,y,z)=(\dot{\gamma}y,0,0)$, and impose a short-range repulsive force $\bm{F}_{\text{ext}}$ between the particles to model excluded-volume interactions. Eq. (3) is then solved using FSD [30] for $N=1024$ at different Pe and $\phi$; see the supplemental material (SM)  [^1] See Supplemental Material at [URL will be inserted by publisher] for a detailed description of our simulation method, setup and verification, which includes Refs. [50, 51, 52, 53]. for further details.
Apolar active suspensions—Recently, we showed that suspensions of shakers, pulling or pushing, could undergo activity-induced diffusion without any external force [31]. The effective diffusion coefficient was found to be $D_{0}=(a^{2}/\tau_{a})g(\phi)$, where $\tau_{a}\equiv a/|B_{2}|$ is the activity time scale, and $g(\phi)\sim\phi^{1/5}$ for $\phi\lessapprox 10\%$ and reduces sharply with $\phi$ beyond. Furthermore, the short-time dynamics was ballistic and could last over tens of $\tau_{a}$; c.f. Fig. 1(b). This persistent motion, therein due to HIs between the active particles, is the main reason that active fluids display an enhanced diffusion [3, 6, 31].
Figure 2: Short-time dynamics and phase diagram. (a) Root-mean-square particle speeds under shear, normalized by their values without shear at all Pe and $\phi$. (b) Normalized velocity autocorrelation for passive particles at different $\phi$ (markers denote $\Phi_{v}(\dot{\gamma}\tau_{d})=-10^{-2}$) and for pulling shakers under different Pe at $\phi=10\%$ (inset). (c) Phase diagram of all simulation results for shakers, divided by Pe $=1$ and 100 into three diffusive regimes with different scalings.
[…]
We extract the diffusion coefficient, $D$, by fitting the long-time MSD according to $\langle\Delta y^{2}(t)+\Delta z^{2}(t)\rangle=4D(t-t_{0})$, where $t_{0}$ is generally nonzero due to nondiffusive dynamics at short times. Fig. 1(c) shows the relative diffusivity, $D/D_{0}$, across four decades of Pe. Clearly, $D/D_{0}$ is nonmonotonic at $\phi=10\%$, with a minimum of about 0.4 reached at Pe $\approx 100$, consistent with the observation in Fig. 1(b). At higher concentrations, the nonmonotonic behavior gradually disappears, and the scaling between $D/D_{0}$ and Pe depends on $\phi$. However, in all cases, we observe $D/D_{0}\approx 1$ for Pe $\lessapprox 1$ and $D/D_{0}\propto$ Pe for Pe $\gtrapprox 100$. The same scalings apply to both pullers and pushers, as is the case without shear [31], implying similar but still irreversible macroscopic dynamics under time reversal.
To explain the nonmonotonic diffusion, we examine the short-time dynamics characterized by the instantaneous particle velocities, projected onto the $yz$-plane as in the MSD calculation. Fig. 2(a) shows the root-mean-square particle speeds under shear, $U_{rms}$, normalized by their values without shear, $U_{rms,0}$, for pullers and pushers at different $\phi$. Remarkably, all data collapse onto a single curve, showing universal scalings with $U_{rms}/U_{rms,0}\approx 1$ for Pe $\lessapprox 1$ and $U_{rms}/U_{rms,0}\propto$ Pe for Pe $\gtrapprox 10$. The absence of any $\phi$-dependence suggests that the relative change of the short-time dynamics is controlled by shear alone, despite the underlying activity-induced dynamics, $U_{rms,0}$ and $D_{0}$, depend nonmonotonically on $\phi$ [31].
[…]
We now explain the observed diffusion. Fig. 2(c) shows the phase diagram of all shaker simulations controlled by two parameters, $\phi$ and $\dot{\gamma}\tau_{a}=g(\phi)$Pe. As discussed, there are three diffusive regimes separated by Pe $\approx 1$ and 100, with different scalings between $D/D_{0}$ and Pe in each. Specifically, Pe $<1$ corresponds to $a^{2}/D_{0}<\dot{\gamma}^{-1}$, i.e., the activity-induced diffusion occurs faster than the shear-induced, thus we observe the former, $D=D_{0}$. When Pe $>1$, both activity and shear are at play, resulting in $\phi$-dependent scalings, including a reduced diffusion at low $\phi$ due to the shear-induced velocity anticorrelation. However, at sufficiently large Pe, the dynamics must be controlled by shear, $D\propto\dot{\gamma}a^{2}$, leading to $D/D_{0}\propto$ Pe. Here, the threshold Pe corresponds to $\tau_{d}=\tau_{a}$, where $\tau_{d}$ is the velocity decorrelation time, i.e., the time $\Phi_{v}(t)$ returns to zero from below, c.f. Fig. 2(b). When $\tau_{d}<\tau_{a}$, diffusion occurs due to shear because even the shortest activity time scale ($\tau_{a}$) is longer than the diffusive time of shear ($\tau_{d}$). Indeed, Fig. 2(c) shows that $\tau_{a}=\tau_{d}$ approximates Pe $=100$, which confirms our analysis.
ABPs—One may wonder whether the nonmonotonic diffusion observed here is a general feature of fluids with persistent internal motion. After all, if it is observable in apolar active suspensions, the effect may be stronger for polar particles with a larger persistence. To test this hypothesis, we perform additional simulations of ABPs, a widely studied model [41] for self-propelled particles, e.g., self-phoretic colloids [42]. Specifically, we simulate neutral squirmers with $B_{1}>0$ and $B_{2}=0$, and impose a rotational diffusion, $d_{r}$, to effect angular noise in the particle dynamics, $\langle\bm{\Omega}_{\text{s}}(t)\cdot\bm{\Omega}_{\text{s}}(t^{\prime})\rangle=2d_{r}\delta(t-t^{\prime})$, where $\delta(t)$ is the Dirac delta. Translational noise is neglected as done typically [43]. Without shear, our ABPs self-propel with a constant speed, $U_{\text{s}}=\frac{2}{3}B_{1}$, unless they collide. The underlying diffusivity is $D_{0}=\frac{1}{3}U_{\text{s}}^{2}\tau_{r}$, where $\tau_{r}=\frac{1}{2}d_{r}^{-1}$, in the dilute regime (see SM [39]), and the persistence length, $\ell_{p}\equiv U_{\text{s}}\tau_{r}$, is tunable.
Figure 3: Diffusivity of ABPs with different persistence length ($\ell_{p}$) at $\phi=10\%$. The main plot shows the relative diffusivity in the $y$ direction; the inset in $z$.
Fig. 3 shows the relative diffusivities, $D_{yy}/D_{0}$ and $D_{zz}/D_{0}$, in $y$ and $z$ directions, separately, for ABPs with different $\ell_{p}$ at $\phi=10\%$ under shear. The diffusivities are extracted in the same way as for shakers. For $D_{yy}/D_{0}$, we observe the expected effect of $\ell_{p}$, i.e., a larger persistence leads to a larger drop in diffusivity, which can be more than 90% if $\ell_{p}/a>3$ (as a reference, $\ell_{p}/a\approx 1.4$ for shakers at $\phi=10\%$ [31]), and the nonmonotonic effect disappears when $\ell_{p}/a\ll 1$. Conversely, $D_{zz}/D_{0}$ always increases with Pe, and the data at different $\ell_{p}$ collapse, scaling as SID. We observe these contrasting behaviors because ABPs hardly interact. As shear rotates all particles around the $z$ axis with a constant speed $\Omega_{\infty}=\dot{\gamma}/2$, ABPs oscillate in the $xy$-plane but remain persistent along $z$ (manifest in the MSDs, see SM [39]). Consequently, $D_{zz}/D_{0}$ can decouple from $D_{yy}/D_{0}$ over a large range of Pe, leading to highly anisotropic diffusion.
The above results depict a simple physical picture of the interplay between shear and activity. When shearing a fluid with an underlying persistence, arising from either activity-induced HIs (as in the case of shakers), self-propulsion (ABPs), or a combination of the two (squirmers in general), the rotational effect of the shear reduces the persistent motion in the shear plane. In the dilute regime where SID is weak, reduced persistence leads to reduced diffusion, as exemplified by a single ABP under shear [44, 45]. However, as the shear rate increases, particles collide more often due to stronger advection, eventually resulting in increased diffusion. The exact behavior depends on the propulsion and interaction mechanisms; nevertheless, the competition between external rotation, internal persistence, and SID determines the diffusive dynamics of active fluids in general.
Figure 4: Diffusivity of passive colloidal suspensions at different $\phi$. The inset shows the MSDs at $\phi=10\%$ under different Pe, where $\tau_{B}\equiv a^{2}/\big{(}k_{\text{B}}T\mu(0)\big{)}$.
Passive colloidal suspensions—The dynamics of ABPs suggests that a persistent internal motion is necessary for observing a nonmonotonic diffusion under shear. To prove this, we consider the contraposition, i.e., passive colloids without any persistence. The underlying diffusivity is $D_{0}=k_{\text{B}}T\mu(\phi)$, where $k_{\text{B}}$ is the Boltzmann’s constant, $T$ the absolute temperature, and $\mu(\phi)$ a decreasing function of $\phi$ [46]. Fig. 4 shows the relative diffusivity, $D/D_{0}$, against Pe for $\phi$ from 10% to 40%, as well as the MSDs under different Pe at $\phi=10\%$. Clearly, shear tends to enhance the diffusion at all $\phi$, despite the subdiffusive dynamics appearing transiently at higher Pe (see Fig. 4 inset and SM [39]). These observations thus verify our hypothesis, highlighting a fundamental difference between active and passive fluids under shear.
Concluding remarks—We have presented a numerical study on the dynamics of sheared particle suspensions using rigorous hydrodynamic simulations. The central finding is that, while shear always enhances the dynamics of passive particles, it can suppress the diffusion in active suspensions by reducing the persistent motion of active particles. This interplay between shear and activity is more pronounced in dilute suspensions, where the activity-induced diffusivity is larger than the shear-induced, and it does not depend on the origin of the persistent motion, being either an intrinsic property of the particles or an emergent phenomenon due to HIs. Therefore, we expect the observed coupling to be a general feature of fluids endowed with an underlying microstructure and internal persistence. Finally, there is currently a growing interest in the relation between the dynamics and rheology of active fluids, which has great practical implications but is still not well understood [45, 47, 48, 49]. Our study may shed some light on this problem.
Acknowledgments—We are grateful to B. Chen for code development, and thank E. Kanso and T. Zhou for discussions. Z.G. acknowledges support from the Swedish Research Council under Grant No. 2021-06669VR.